the no of combination 16 things in which 8 things are alike and the rest different taken 8 at a time is??

Grade:11

3 Answers

Garvita Singhvi
8 years ago
allike-A diffrent-D
alll A -1
7A 1D-8
6A 2D- 8C2=28
5A 3D -8C3= 56
4A 4D -8C4= 70
3A 5D-8C5=56
2A 6D-8C6=28
all D-9C8=9
TOTAL=1+8+28+56+70+56+28+9=265
sourav
6 years ago
8c8+8c7+8c6+8c5+8c4+8c3+8c2+8c1+9c8=256At last we used 9c8 because there can be 9 different items at a time
Parth
5 years ago
 
A -1
7A 1D-8
6A 2D- 8C2=28
5A 3D -8C3= 56
4A 4D -8C4= 70
3A 5D-8C5=56
2A 6D-8C6=28
Answer   256
all D-9C8=9
Answer   256